OpenStudy (anonymous):

Show that 9+90+900+...+9x19^n-1=10^n -1 (Arithmetic and geometric series and their sums)

OpenStudy (anonymous):

what about 19^n-1, i think it should be 10^n-1 9+90+900+...n terms =9(1+10+100+...n terms) ...(1) \[1+10+100+... n terms.\] it is a G.P. first term a=1 common ratio r=10 \[S=a \frac{ r ^{n}-1 }{r-1 }\] find S and put in (1)

OpenStudy (anonymous):

let S = 9 + 90 + 900 + ... S = 9(10)^0 + 9(10)^1 + 9(10)^2 + .. + 9(10)^(n-1) then 10S = 9(10)^1 + 9(10)^2 + 9(10)^3 + .. + 9(10)^n subtract the two equations: 10S - S = 9(10)^n - 9 9S = 9(10)^n - 9 S = 10^n - 1